# Reconstruction Alignment Improves Unified Multimodal Models

### v1

@@ Abstract
Unified multimodal models (UMMs) unify visual understanding and generation within a single architecture. However, conventional training relies on image–text pairs (or sequences) whose captions are typically sparse and miss fine-grained visual details—even when they use hundreds of words to describe a simple image. We introduce Reconstruction Alignment (RecA), a resource-efficient post-training method that leverages visual understanding encoder embeddings as dense “text prompts,” providing rich supervision without captions. Concretely, RecA conditions a UMM on its own visual understanding embeddings and optimizes it to reconstruct the input image with a self-supervised reconstruction loss, thereby realigning understanding and generation. Despite its simplicity, RecA is broadly applicable: across autoregressive, masked-autoregressive, and diffusion-based UMMs, it consistently improves generation and editing fidelity. With only 27 GPU-hours, post-training with RecA substantially improves image generation performance on GenEval (0.73→0.90) and DPGBench (80.93→88.15), while also boosting editing benchmarks (ImgEdit 3.38→3.75, GEdit 6.94→7.25). Notably, RecA surpasses much larger open-source models and applies broadly across diverse UMM architectures, establishing it as an efficient and general post-training alignment strategy for UMMs.
@@ 1 Introduction
Building on the success of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2024a), researchers have developed Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2023; 2025; Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024; Zhu et al., 2025) with strong visual understanding performance. Recently, unified multimodal models (UMMs), or Omni Models, have been proposed to both understand and generate across modalities—reading and writing visual content and text within a single architecture (Team, 2024; Zhou et al., 2025; Tong et al., 2024; Ge et al., 2024; Wu et al., 2024b; Chen et al., 2025c; Pan et al., 2025a; OpenAI, 2024; Li et al., 2025). The academic community envisions that a unified framework can inherit the reasoning and world knowledge of LLMs while extending them to content generation.
@@
Building on these insights, we propose RecA, a resource-efficient post-training strategy. The core idea is simple: condition UMMs on their own understanding vision encoder embeddings—dense “visual prompts” that encode layout, color, and attributes beyond what captions capture—and train them to reconstruct the image. This semantic reconstruction provides richer supervision without additional labels. Despite its simplicity, RecA yields substantial improvements. With only 27 A100 GPU hours, a 1.5B-parameter UMM post-trained with RecA surpasses GPT-4o and significantly larger open-source models, achieving state-of-the-art results on GenEval (0.86) and DPGBench (87.21). Importantly, these gains are obtained without any GPT-4o-Image distillation data or reinforcement learning (OpenAI, 2024; Chen et al., 2025a; b), in contrast to prior work. Moreover, when post-trained with GPT-4o data, RecA further improves to GenEval (0.90) and DPGBench (88.15), substantially surpassing previous methods that use much larger model sizes. RecA also boosts image editing quality, raising ImgEdit from 3.38→3.75 and GEdit from 6.94→7.25. Moreover, RecA applies broadly across UMM families—including Show-o (Xie et al., 2025b) (AR), Harmon (Wu et al., 2025c) (AR+MAR), OpenUni (Wu et al., 2025b) (AR+Diffusion), and BAGEL (Deng et al., 2025) (AR+Diffusion)—highlighting its generality.
@@
  • Performance: We demonstrate strong empirical gains, with a 1.5B-parameter model surpassing GPT-4o and larger open-source models using only 27 A100 GPU hours, significantly outperforming prior state of the art without distillation or RL.
Figure 4: Overview of the semantic reconstruction alignment (RecA) pipeline. A visual understanding encoder (e.g., CLIP or DINO) extracts semantic features from the input image, which are fused with template text embeddings and passed to a Unified Multimodal Model (UMM) to regenerate the image. The UMM is optimized with a self-supervised loss (diffusion or cross-entropy) between the original and reconstructed images or image latents. At inference time, RecA requires no additional inputs, operating as a standard UMM.
@@ 2.2 RecA Training Paradigm
[…]
@@ 2.3 Difference Between RecA and Previous Works
RecA introduces semantic-level image reconstruction as a native post-training objective for UMMs, directly leveraging visual understanding priors to improve both generation and editing, without auxiliary modules or text-image supervision. In contrast, previous works integrate reconstruction in different ways: (I) diffusion-supervised enhancement (e.g., DIVA, ViLex) (Wang et al., 2024b; 2025c; Ma et al., 2025a) leverages pretrained diffusion models to regularize vision encoders and improve understanding; (II) reconstruction from hidden states (e.g., ROSS (Wang et al., 2024a; 2025b)) adds lightweight decoders to regenerate input images from intermediate embeddings, thereby regularizes the model to preserve fine-grained details for VLMs; (III) representation alignment (e.g., REPA (Yu et al., 2024)) introduces additional alignment modules that map noisy hidden states from the VAE in denoising networks to clean image representations obtained from external, pretrained visual encoders; and (IV) reconstruction as a prior (e.g., Lumos (Ma et al., 2025b)) adds additional DINO features (Caron et al., 2021) into the attention blocks of the diffusion model, which is further trained on large-scale text-to-image data. Our method is fundamentally different in terms of the methodology, architecture, motivation and task.
@@ 3.1 Experiment Setup
  • Harmon (MAR) (Wu et al., 2025c) adopts masked autoregressive generation, with MAE (Li et al., 2024) / VAE (Kingma & Welling, 2013) as the visual understanding / generation encoder, evaluated at 0.5B and 1.5B parameter scales.
  • OpenUni (AR+Diff) (Wu et al., 2025b) integrates autoregressive and diffusion generation, with InternVL3 (Zhu et al., 2025) / VAE (Kingma & Welling, 2013) as the visual understanding / generation encoder. It serves as an open-source counterpart of MetaQueries (Pan et al., 2025b); we evaluate both 1.6B and 3.6B variants at $\times512$ resolution without GPT-4o-Image distillation.
@@
Together, these models cover the main families of UMM architectures: autoregressive (AR), masked autoregressive (MAR), and autoregressive+diffusion (AR+Diff).
Evaluation details. We evaluate text-to-image generation on GenEval (Ghosh et al., 2023) and DPGBench (Hu et al., 2024), and image editing on ImgEdit (Ye et al., 2025) and GEdit-Bench-EN (Liu et al., 2025). Our baselines include both generation-only models (e.g., SDXL (Podell et al., 2023), DALL-E 3 (Betker et al., )) and unified multimodal models (e.g., Show-o (Xie et al., 2025b), Harmon (Wu et al., 2025c), BAGEL (Deng et al., 2025), GPT-4o-Image (OpenAI, 2024)).
@@ 3.2 RecA Achieves SOTA Performance on Image Generation and Editing
SOTA performance without GPT-4o-Image distillation. As shown in Table 1, post-training Harmon-1.5B model with RecA achieves SOTA results across both generation-only and UMM architectures. Despite using significantly fewer parameters, RecA surpasses recent methods, e.g., Janus-Pro (Chen et al., 2025c), OmniGen2 (Wu et al., 2025a), and BAGEL (Deng et al., 2025).
@@ 3.3 RecA is Effective Across Different UMM Frameworks
Performance on WISE benchmark. WISE (Niu et al., 2025) evaluates text-to-image reasoning through 1,000 knowledge-puzzle prompts requiring implicit factual knowledge (e.g., Einstein’s favorite instrument). As shown in Table 2, RecA yields consistent gains on Harmon and OpenUni, while BAGEL and Show-o show smaller improvements. These results highlight that RecA is particularly effective at strengthening semantic alignment, while deeper reasoning may depend more on advances in the underlying language model. We view WISE as a complementary challenge and an exciting direction for future research (Ye et al., 2024; Tong et al., 2024). Additional results are provided in Appendix’s Sec. D.
@@ 3.4 More Results
Table 5: Comparison of SFT and RecA across post-training datasets. Post-training with RecA consistently outperforms SFT, especially when benchmark-specific data (e.g., GenEval) is excluded. On high-quality images, RecA can even surpass SFT under dense prompts. All results are reported at 5k steps.
@@ 3.5 Empirical Studies of RecA as a Post-Training Method
RecA is more effective for post-training. As shown in Table 5, on MidjourneyV6, SFT yields modest improvements but quickly plateaus at 74.76 on GenEval. Adding denser captions via Midjourney-LLaVA (brivangl, 2024) does not help—GenEval remains nearly unchanged and DPGBench even drops to 80.67. In contrast, replacing SFT with RecA dramatically boosts performance under the same setup, achieving 85.69 GenEval (+10.93) and 87.21 DPGBench (+6.54). This demonstrates that self-supervised semantic reconstruction is a far more effective post-training strategy than supervised fine-tuning. Further analysis of SFT on BLIP3o-60k is provided in Appendix C.
@@ 4.1 Unified Multimodal Models (UMMs)
Unified multimodal model (UMM) is a single, end-to-end-trained backbone capable of multimodal understanding and generation within a single model or system (Zhang et al., 2025b). Recent studies have explored a broad design space for different UMMs: (I) AR. Chameleon and Janus tokenize images and predict them autoregressively in a next-token prediction way (Team, 2024; Qu et al., 2024; Chen et al., 2025d; c; Wu et al., 2024a), and Show-o (Xie et al., 2025b) introduces a discrete-diffusion schedule to improve the token prediction process. (II) AR + Diffusion. Models such as Transfusion attach a diffusion (or flow-matching) head to the shared transformer (Zhou et al., 2025; Deng et al., 2025; Ma et al., 2024; Shi et al., 2024; Zhang et al., 2025a; Xie et al., 2025c). Some UMMs keep a powerful pretrained MLLM frozen for reasoning and route its intermediate features with learnable queries (Pan et al., 2025b; Chen et al., 2025a; Wu et al., 2025b; Lin et al., 2025) to an external image generator. (III) AR + MAR. Masked-autoregressive (MAR)  (Li et al., 2024) is a novel autoregressive image generation paradigm without any vector quantization, which has been adopted by models like Harmon (Wu et al., 2025c; Fan et al., 2025; Yang et al., 2024b).
@@ 4.2 Post-training Strategies for UMMs
Recent efforts explore different post-training techniques for enhancing the generation capability of UMMs. (I) Chain-of-Thought (CoT) or test-time verification let the model do reasoning before the generation or verify generated image step-by-step, but they depend on external models and do not improve the UMM’s native generation capability.  (Guo et al., 2025; Wang et al., 2025d; Fang et al., 2025; Tian et al., 2025). (II) Reinforcement learning. Methods such as DPO and GRPO optimize generation policies with human or automatic preference signals, but require curated paired data and carefully tuned advantage functions.  (Guo et al., 2025; Mao et al., 2025; Tian et al., 2025; Jiang et al., 2025). (III) High-quality synthetic data. BLIP3o distils 60,000 GPT-4o-Image generated text-image pair for GenEval, dense-prompt, pose, and landmark scenarios, while ShareGPT-4o-Image contributes rich editing data (OpenAI, 2024; Chen et al., 2025a; b; Wang et al., 2025e). Supervised fine-tuning (SFT) on these sets lifts benchmark scores appreciably, yet the approach needs heavy data-generation cost and risks distribution shift.
@@ 5 Conclusion
In this work, we propose RecA—a lightweight post-training paradigm that replaces sparse text-to-image supervision with dense features from the model’s own visual understanding encoder. RecA requires no extra caption data yet improves image generation and editing across architectures: our RecA-1.5B reaches 0.86 on GenEval and 87.21 on DPGBench, and a sequential two-staged strategy pushes them to 0.90 and 88.15, setting new state-of-the-art. We discuss the broader implications, limitations and future directions of our work in Appendix C.
@@ Acknowledgments
We sincerely thank Ishan Misra, Xingyi Zhou, Haoqi Fan, Mannat Singh, Dewei Zhou, Haotian Tang, Yulun Wu, Baifeng Shi, Jiaxin Ge, Yuwei Niu, Xun Wang, Jiuhai Chen for their insightful discussions and valuable feedback on our paper. We especially thank Size Wu and Yang Ye for their assistance in reproducing the baseline results.
@@ Appendix B Experimental Setup
Evaluation Details. Following previous work, we evaluate text-to-image generation capabilities using GenEval (Ghosh et al., 2023) and DPGBench (Hu et al., 2024). Our baselines include: (I) Generation-only models: SD3-Medium (Esser et al., 2024), SDXL (Podell et al., 2023), SANA-1.5 (Xie et al., 2025a), Emu3-Gen (Wang et al., 2024c), FLUX-dev (BlackForest, 2024), Playground-v3 (Liu et al., 2024a), and DALL-E 3 (Betker et al., ). (II) Unified multimodal models: Show-o (Xie et al., 2025b), Harmon (Wu et al., 2025c), Janus-pro (Chen et al., 2025c), OmniGen2 (Wu et al., 2025a), BLIP3-o (Chen et al., 2025a), BAGEL (Deng et al., 2025), Show-o2 (Xie et al., 2025c), UniWorld-V1 (Lin et al., 2025), Ovis-U1 (Wang et al., 2025a), and GPT-4o-Image (OpenAI, 2024). For enhanced statistical reliability, we re-evaluate some models using 12 random seeds (tripling the standard 4 seeds).
@@
Training Data. Due to limited availability of UMM training data, we use high-quality open-source MidjourneyV6 dataset (CortexLM, 2024) for RecA. To preserve visual understanding capabilities of UMMs with shared parameters for both understanding and generation (Show-o and Harmon), we also incorporate LLaVA Mix-665K (Liu et al., 2024b). For BAGEL, which shows sensitivity to image distribution, we use 8,000 $1024\times 1024$ FLUX-generated samples from Text-to-Image-2M (jackyhate, 2024).
@@
$\lambda_{\text{i2t}}$ 1.0 1.0 1.0 0.0 0.0 0.0
$\lambda_{\text{{RecA}}}$ 1.0 1.0 1.0 1.0 1.0 1.0

### v2

@@ Abstract
Unified multimodal models (UMMs) unify visual understanding and generation within a single architecture. However, conventional training relies on image–text pairs (or sequences) whose captions are typically sparse and miss fine-grained visual details—even when they use hundreds of words to describe a simple image. We introduce Reconstruction Alignment (RecA), a resource-efficient post-training method that leverages visual understanding encoder embeddings as dense “text prompts,” providing rich supervision without captions. Concretely, RecA conditions a UMM on its own visual understanding embeddings and optimizes it to reconstruct the input image with a self-supervised reconstruction loss, thereby realigning understanding and generation. Despite its simplicity, RecA is broadly applicable: across autoregressive, masked-autoregressive, and diffusion-based UMMs, it consistently improves generation and editing fidelity. With only 27 GPU-hours, post-training with RecA substantially improves image generation performance on GenEval (0.73→0.90) and DPGBench (80.93→88.15), while also boosting editing benchmarks (ImgEdit 3.38→3.75, GEdit 6.94→7.27). Notably, RecA surpasses much larger open-source models and applies broadly across diverse UMM architectures, establishing it as an efficient and general post-training alignment strategy for UMMs.
@@ 1 Introduction
Building on the success of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2024a), researchers have developed Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2023; 2025; Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024; Zhu et al., 2025) with strong visual understanding performance. Recently, unified multimodal models (UMMs), or Omni Models, have been proposed to both understand and generate across modalities—reading and writing visual content and text within a single architecture (Tong et al., 2024a; Ge et al., 2024; Wu et al., 2024b; Chen et al., 2025c; Pan et al., 2025a; OpenAI, 2024; Li et al., 2025; Wu et al., 2025a). The academic community envisions that a unified framework can inherit the reasoning and world knowledge of LLMs while extending them to content generation.
@@
Building on these insights, we propose RecA, a resource-efficient post-training strategy. The core idea is simple: condition UMMs on their own understanding vision encoder embeddings—dense “visual prompts” that encode layout, color, and attributes beyond what captions capture—and train them to reconstruct the image. This semantic reconstruction provides richer supervision without additional labels. Despite its simplicity, RecA yields substantial improvements. With only 27 A100 GPU hours, a 1.5B-parameter UMM post-trained with RecA surpasses GPT-4o and significantly larger open-source models, achieving state-of-the-art results on GenEval (0.86) and DPGBench (87.21). Importantly, these gains are obtained without any GPT-4o-Image distillation data or reinforcement learning (OpenAI, 2024; Chen et al., 2025a; b), in contrast to prior work. Moreover, when post-trained with GPT-4o data, RecA further improves to GenEval (0.90) and DPGBench (88.15), substantially surpassing previous methods that use much larger model sizes. RecA also boosts image editing quality, raising ImgEdit from 3.38→3.75 and GEdit from 6.94→7.27. Moreover, RecA applies broadly across UMM families—including Show-o (Xie et al., 2025b) (AR), Harmon (Wu et al., 2025d) (AR+MAR), OpenUni (Wu et al., 2025c) (AR+Diffusion), and BAGEL (Deng et al., 2025) (AR+Diffusion)—highlighting its generality.
@@
  • Performance: We demonstrate strong empirical gains, with a 1.5B-parameter model surpassing GPT-4o and larger open-source models using only 27 A100 GPU hours, significantly outperforming prior state of the art without distillation or reinforcement learning (RL).
Figure 4: Overview of the semantic reconstruction alignment (RecA) pipeline. A visual understanding encoder (e.g., CLIP or DINO) extracts semantic features from the input image, which are fused with template text embeddings and passed to a Unified Multimodal Model (UMM) to regenerate the image. The UMM is optimized with a self-supervised loss (diffusion or cross-entropy) between the original and reconstructed images or image latents. We freeze the visual understanding encoder, except in cases where the UMM employs the same encoder for both understanding and generation (e.g., Harmon, which uses an MAE encoder for both tasks). At inference time, RecA requires no additional inputs, operating as a standard UMM.
@@ 2.2 RecA Training Paradigm
Training configuration. We set $\lambda_{\text{{RecA}}}=1,\lambda_{\text{t2i}}=0$. For UMMs that share parameters between understanding and generation, we set $\lambda_{\text{i2t}}=1$ to preserve image-to-text capability; otherwise, we set $\lambda_{\text{i2t}}=0$ and freeze the understanding component. We freeze the visual understanding encoder in most cases, except when the UMM employs the same encoder for both tasks. Input images are resized to the minimum resolution accepted by the visual encoder.
@@ 2.3 Difference Between RecA and Previous Works
RecA introduces semantic-level image reconstruction as a native post-training objective for UMMs, directly leveraging visual understanding priors to improve both generation and editing, without auxiliary modules or text-image supervision. In contrast, previous works integrate reconstruction in different ways: (I) diffusion-supervised enhancement (e.g., DEEM, DIVA, ViLex) (Luo et al., 2024; Wang et al., 2024b; 2025d; Ma et al., 2025a) leverages pretrained diffusion models to regularize vision encoders and improve understanding; (II) reconstruction from hidden states (e.g., ROSS) (Wang et al., 2024a; 2025b)) adds lightweight decoders to regenerate input images from intermediate embeddings, thereby regularizes the model to preserve fine-grained details for VLMs; (III) representation alignment (e.g., REPA) (Yu et al., 2024) introduces additional alignment modules that map noisy hidden states from the VAE in denoising networks to clean image representations obtained from external, pretrained visual encoders; and (IV) reconstruction as a prior (e.g., Lumos) (Ma et al., 2025b) adds additional DINO features (Caron et al., 2021) into the attention blocks of the diffusion model, which is further trained on large-scale text-to-image data. Our method is fundamentally different in terms of the methodology, architecture, motivation and task.
@@ 3 Experiments
  • Robustness: RecA consistently improves generation capabilities across diverse datasets and benchmarks, indicating that the gains are not from training-data memorization. (Table 5)
@@ 3.1 Experiment Setup
  • Harmon (AR+MAR) (Wu et al., 2025d) adopts masked autoregressive generation, with MAE (Li et al., 2024) / MAE+VAE (Kingma & Welling, 2013) as the visual understanding / generation encoder, evaluated at 0.5B and 1.5B parameter scales.
  • OpenUni (AR+Diff) (Wu et al., 2025c) integrates autoregressive and diffusion generation, with InternVL3 (Zhu et al., 2025) / VAE (Kingma & Welling, 2013) as the visual understanding / generation encoder. It serves as an open-source counterpart of MetaQueries (Pan et al., 2025b); we evaluate both 1.6B and 3.6B variants at $\times512$ resolution without GPT-4o-Image distillation.
@@
Together, these models cover the main families of UMM architectures: autoregressive (AR), autoregressive+masked autoregressive (AR+MAR), and autoregressive+diffusion (AR+Diff).
Evaluation details. We evaluate text-to-image generation on GenEval (Ghosh et al., 2023) and DPGBench (Hu et al., 2024), and image editing on ImgEdit (Ye et al., 2025) and GEdit-Bench-EN (Liu et al., 2025). Our baselines include both generation-only models (e.g., SDXL (Podell et al., 2023), DALL-E 3 (Betker et al., )) and unified multimodal models (e.g., Show-o (Xie et al., 2025b), Harmon (Wu et al., 2025d), BAGEL (Deng et al., 2025), GPT-4o-Image (OpenAI, 2024)).
@@ 3.2 RecA Achieves SOTA Performance on Image Generation and Editing
SOTA performance without GPT-4o-Image distillation. As shown in Table 1, post-training Harmon-1.5B model with RecA achieves SOTA results across both generation-only and UMM architectures. Despite using significantly fewer parameters, RecA surpasses recent methods, e.g., Janus-Pro (Chen et al., 2025c), OmniGen2 (Wu et al., 2025b), and BAGEL (Deng et al., 2025).
@@ 3.3 RecA is Effective Across Different UMM Frameworks
Performance on WISE benchmark. WISE (Niu et al., 2025) evaluates text-to-image reasoning through 1,000 knowledge-puzzle prompts requiring implicit factual knowledge (e.g., Einstein’s favorite instrument). As shown in Table 2, RecA yields consistent gains on Harmon and OpenUni, while BAGEL and Show-o show smaller improvements. These results highlight that RecA is particularly effective at strengthening semantic alignment, while deeper reasoning may depend more on advances in the underlying language model. We view WISE as a complementary challenge and an exciting direction for future research (Ye et al., 2024; Tong et al., 2024a). Additional results are provided in Appendix’s Sec. D.
@@ 3.4 More Results
Table 5: Comparison of SFT and RecA across post-training datasets. RecA consistently outperforms SFT on GenEval, particularly when template leakage data is excluded (yellow row), demonstrating that the improvements are genuine and not due to memorizing specific data patterns. Under the dense-prompt setting, RecA easily surpasses SFT. All results are reported at 5k steps.
@@ 3.5 Empirical Studies of RecA as a Post-Training Method
RecA demonstrates genuine effectiveness compared to SFT. As shown in Table 5, on MidjourneyV6, SFT achieves 74.76 on GenEval. Adding denser captions via Midjourney-LLaVA (brivangl, 2024) does not help—GenEval remains nearly unchanged and DPGBench even drops to 80.67. In contrast, replacing SFT with RecA dramatically boosts performance under the same setup, achieving 85.69 GenEval (+10.93) and 87.21 DPGBench (+6.54). On the complete BLIP3o-60k dataset, RecA still outperforms SFT on both GenEval and DPGBench. Notably, removing template leakage data leads to a performance drop for SFT, whereas RecA maintains strong performance—clearly demonstrating that RecA provides genuine improvements that are not simply memorizing the training data.
@@ 4.1 Unified Multimodal Models (UMMs)
Unified multimodal model (UMM) is a single, end-to-end-trained backbone capable of multimodal understanding and generation within a single model or system (Zhang et al., 2025b). Recent studies have explored a broad design space for different UMMs: (I) AR. Models like Chameleon tokenize images and predict them autoregressively in a next-token prediction way (Team, 2024; Qu et al., 2024; Chen et al., 2025d; c; Wu et al., 2024a), while Show-o (Xie et al., 2025b) introduces a discrete-diffusion schedule to improve the token prediction process. (II) AR + Diffusion. Models like Transfusion attach a diffusion (or flow-matching) (Ho et al., 2020; Lipman et al., 2023) head to the shared transformer (Zhou et al., 2025; Deng et al., 2025; Ma et al., 2024; Shi et al., 2024; Zhang et al., 2025a; Xie et al., 2025c), while some UMMs keep pretrained MLLMs frozen and route its intermediate features with learnable queries to an external image generator (Pan et al., 2025b; Chen et al., 2025a; Wu et al., 2025c; Lin et al., 2025). (III) AR + MAR. Masked-autoregressive (MAR)  (Li et al., 2024) is a novel autoregressive image generation paradigm without any vector quantization, which has been adopted by models like Harmon (Wu et al., 2025d; Fan et al., 2025; Yang et al., 2024b; Wang et al., 2025c).
@@ 4.2 Post-training Strategies for UMMs
Recent efforts explore different post-training techniques for enhancing the generation capability of UMMs. (I) Chain-of-Thought (Wei et al., 2022) or test-time verification do reasoning before the generation or verify generated image step-by-step, but they depend on external models and do not improve the UMM’s native generation capability.  (Guo et al., 2025; Wang et al., 2025e; Fang et al., 2025; Tian et al., 2025). (II) Reinforcement learning. Methods such as DPO (Rafailov et al., 2023) and GRPO (Shao et al., 2024) optimize generation policies with human or automatic preference signals, but require curated paired data and carefully tuned advantage functions.  (Wei et al., 2025; Guo et al., 2025; Mao et al., 2025; Tian et al., 2025; Jiang et al., 2025). (III) High-quality synthetic data. Recent works (e.g., BLIP3o)  (Chen et al., 2025a; Wang et al., 2025f)) construct large-scale GPT-4o–Image (OpenAI, 2024) generated image–text pairs for supervised fine-tuning, which boost GenEval and editing benchmarks but incur high data-generation cost and potential distribution shift.
@@ 5 Conclusion
In this work, we propose RecA—a lightweight post-training method that replaces sparse text-to-image supervision with dense features from the model’s own visual understanding encoder. RecA requires no extra caption data yet sigificantly improves image generation and editing across architectures. We discuss the broader implications, limitations and future directions of our work in Appendix C.
@@ Acknowledgments
We sincerely thank Ishan Misra, Xingyi Zhou, Haoqi Fan, Mannat Singh, Dewei Zhou, Haotian Tang, Yulun Wu, Baifeng Shi, Jiaxin Ge, Yuwei Niu, Xun Wang, Jiuhai Chen, Weiyang Jin, Xichen Pan for their insightful discussions and valuable feedback on our paper. We especially thank Size Wu and Yang Ye for their assistance in reproducing the baseline results.
@@ Appendix B Experimental Setup
Evaluation Details. Following previous work, we evaluate text-to-image generation capabilities using GenEval (Ghosh et al., 2023) and DPGBench (Hu et al., 2024). Our baselines include: (I) Generation-only models: SD3-Medium (Esser et al., 2024), SDXL (Podell et al., 2023), SANA-1.5 (Xie et al., 2025a), Emu3-Gen (Wang et al., 2024c), FLUX-dev (BlackForest, 2024), Playground-v3 (Liu et al., 2024a), and DALL-E 3 (Betker et al., ). (II) Unified multimodal models: Show-o (Xie et al., 2025b), Harmon (Wu et al., 2025d), Janus-pro (Chen et al., 2025c), OmniGen2 (Wu et al., 2025b), BLIP3-o (Chen et al., 2025a), BAGEL (Deng et al., 2025), Show-o2 (Xie et al., 2025c), UniWorld-V1 (Lin et al., 2025), Ovis-U1 (Wang et al., 2025a), and GPT-4o-Image (OpenAI, 2024). For enhanced statistical reliability, we re-evaluate some models using 12 random seeds (tripling the standard 4 seeds).
@@
Training Data. Due to limited availability of UMM training data, we use high-quality open-source MidjourneyV6 dataset (CortexLM, 2024) for Show-o, Harmon and OpenUni. As we discussed in Sec. 2.2, to preserve visual understanding capabilities of UMMs with shared parameters for both understanding and generation (Show-o and Harmon), we also incorporate LLaVA Mix-665K (Liu et al., 2024b); for BAGEL and OpenUni, we set $\lambda_{\text{i2t}}=0$. Specially For BAGEL, which shows sensitivity to image distribution, we use 8,000 $1024\times 1024$ FLUX-generated samples from Text-to-Image-2M (jackyhate, 2024) (BAGEL is trained on FLUX-generate image).
@@
$\lambda_{\text{i2t}}$ 1.0 1.0 0.0 0.0
$\lambda_{\text{{RecA}}}$ 1.0 1.0 1.0 1.0